OpenStudy (anonymous):

Will Give Metal!! Once Someone Shows Me How To Give One Hahaahah Solve the logarithmic equation. Round to the nearest ten-thousandth if necessary. log 8x^3 = 4 please show all work, thank you :)

OpenStudy (anonymous):

I can't answer that. But to give someone a metal you click 'Best Answer' on the side of their post. Good luck with the question:)

OpenStudy (anonymous):

ohh I see thanks :)

OpenStudy (anonymous):

no problem!

OpenStudy (anonymous):

does anyone know how to do logarithms?

OpenStudy (anonymous):

I found the result, but I'm not sure this is correct! You have the answer?

OpenStudy (anonymous):

no i dont :( i don't understand these at all. What was your steps?

OpenStudy (anonymous):

I did that way: log8x³ = 4 3log8x = 4 3log2³x = 4 3(3)log2x = 4 -> log 2x (approximately 0.30! this value I already knew!) 9.(0,30x) = 4 2,7x = 4 x = 4/2,7 x = 1,48 I do not know if it is correct! : S

OpenStudy (anonymous):

whats the coma for in the answer?

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20130107174304AAEsgfd check out how this person did the problem. the final step they gave, can you see if you understand how to do tht? thank you!!

OpenStudy (anonymous):

x = (10^(4/3))/8 this was the last step.. how do i do that??

OpenStudy (anonymous):

I think this way would not be very correct! 'd Have to use calculator, he did almost the same thing I did!

OpenStudy (anonymous):

A base e solution is attached.

OpenStudy (anonymous):

wow that's confusing haha

OpenStudy (anonymous):

yes :/

OpenStudy (anonymous):

please, in English @robtobey ? Haha its okay, not everyone are geniuses :)